# Relative Camera Pose Regression

### v1

@@ Abstract
Pre-training techniques have greatly advanced computer vision, with CroCo’s cross-view completion approach yielding impressive results in tasks like 3D reconstruction and pose regression. However, this method requires substantial overlap between training pairs, limiting its effectiveness. We introduce Alligat0R, a novel pre-training approach that reformulates cross-view learning as a co-visibility segmentation task. Our method predicts whether each pixel in one image is co-visible in the second image, occluded, or outside the field of view (FOV), enabling the use of image pairs with any degree of overlap and providing interpretable predictions. To support this, we present Cub3, a large-scale dataset with 2.5 million image pairs and dense co-visibility annotations derived from the nuScenes dataset. This dataset includes diverse scenarios with varying degrees of overlap. The experiments show that Alligat0R significantly outperforms CroCo in relative pose regression, especially in scenarios with limited overlap. Alligat0R and Cub3 will be made publicly available.
Figure 1: We introduce Alligat0R, a pretraining method for binocular vision robust to wide baselines. Alligat0R explicitly segments pixels as co-visible, occluded, or outside field-of-view, overcoming the fundamental limitation of CroCo [45, 46] which attempts to reconstruct potentially unobservable regions. We demonstrate Alligat0R on relative pose regression and show that it allows us to significantly outperform the state-of-the-art, especially on pairs with very low overlap.
@@ 1 Introduction
Pre-training techniques have revolutionized computer vision by enabling large models to learn rich representations [3, 28, 5, 17, 16, 1, 53]. In 3D computer vision, CroCo [45, 46] pioneered cross-view completion as a pretext task, where one view is partially masked and reconstructed using visible portions along with a second reference view of the same scene (see Fig. 1 left). This approach allows to learn powerful 3D features and has made possible impressive results on downstream tasks such as 3D reconstruction [43, 12, 47, 40], 3D pose regression [7], camera calibration [24], 4D reconstruction [50, 21, 42] and Gaussian splatting [33, 49].
Despite its success, the cross-view completion pretext task has a fundamental limitation: pretraining is only effective when training pairs have substantial overlap. For instance, Croco [45, 46] relies on pairs with at least 50% overlap. We are particularly interested in relative pose regression between two images, and using this pretext task impairs the performance of pose regression on pairs with limited overlap, as our results show.
In this paper, we present Alligat0R (see Fig. 1 right), a novel pre-training approach that reformulates the cross-view learning objective as a co-visibility segmentation task. Instead of reconstructing masked regions, our method explicitly predicts whether each pixel in one image is: (1) co-visible in the second image, (2) occluded, or (3) outside the field of view (FOV). This formulation offers several advantages: it can use image pairs with any degree of overlap, it provides interpretable predictions that reveal the model’s geometric understanding, and it aligns more directly with the correspondence reasoning required in downstream binocular vision tasks.
To enable our approach, we introduce Cub3, a large-scale dataset comprising 2.5 million image pairs with dense co-visibility annotations derived from the nuScenes dataset. This dataset includes challenging scenarios with varying degrees of overlap between views, providing a more diverse training signal than previous approaches.
One might notice that Alligat0R is not strictly self-supervised, unlike some earlier pre-training methods, as it relies on co-visibility annotations. However, the same can be said for CroCo, which requires filtering image pairs to retain only those with at least 50% overlap. The creation of our co-visibility annotations is more complex than for CroCo’s, however the process is fully automated. This process is very similar to the one used in RUBIK [26] and relies on registered video sequences and depth predictions.
@@
    We introduce co-visibility segmentation as a novel pre-training objective for binocular vision tasks, replacing the cross-completion-based approach of prior work while maintaining the same architecture.
@@
    We demonstrate that co-visibility segmentation-based pretraining significantly benefits from its ability to learn from challenging image pairs, i.e. from any overlap.
@@
    We create and release Cub3, a large-scale dataset with dense co-visibility annotations derived from the nuScenes dataset.
@@
    We show that Alligat0R significantly outperforms CroCo pre-training when fine-tuned on the relative pose regression task, particularly for challenging scenarios with limited overlap between views.
  5. 5.
    We demonstrate the effectiveness of our learning based on geometric difficulty and provide insights into the model’s cross-view reasoning capabilities through interpretable segmentation outputs.
Our experiments demonstrate that explicitly learning to understand co-visibility relationships between image pairs leads to more robust and transferable features for relative pose regression compared to reconstruction-based approaches.
@@ 2 Related Work
To the best of our knowledge, CroCo [45] pioneered the extension of masked image modeling [18] to image pairs, introducing cross-view completion. CroCo v2 [46] applied this cross-view completion framework to large amounts of data. P-Match [54] introduced a variation where both images are partially masked to pre-train an image matching model. Another variant, masked appearance transfer, was proposed in [51] for object tracking, with a similar approach in [34]. Let us highlight that the foundational models DUSt3R [43] and MASt3R [24] are built on CroCo. While all these works rely on cross-view completion, we depart from this framework and propose a novel pretraining objective based on co-visibility segmentation.
Considering co-visibility across image pairs.
Ground-truth co-visibility masks are widely used in image matching methods [30, 10, 25, 52, 35, 44, 11, 4, 41, 37, 9, 13, 15, 23, 14, 20, 36, 31] to compute overlaps and select training pairs. Several approaches [31, 11, 38, 9, 14] also leverage these masks to learn matchability scores, which are used at test time to select correspondences. Additionally, [19] jointly learns co-visibility and relative planar pose between 360 panoramas. To the best of our knowledge, our work is the first to learn visual representations from co-visibility segmentation.
@@
Reloc3R [7] recently fine-tuned the foundational model DUSt3R [43], which is based on CroCo, on large-scale data for relative pose regression. In our experiments, we adopt a similar strategy to evaluate our novel pretraining method against CroCo. However, we focus on the autonomous driving scenario with metric relative translation, whereas Reloc3R predicts only the direction of the relative translation.
@@ 3 Method
In this section, we describe our novel pre-training approach, Alligat0R, which replaces the cross-view completion objective used in CroCo [45, 46] by a segmentation task. We first outline the overall architecture, which remains largely similar to CroCo, and then detail our co-visibility segmentation pre-training objective. Finally, we explain how our model can be fine-tuned for downstream tasks such as relative pose regression.
@@ 3.1 Architecture Overview
Figure 2: Overview of Alligat0R. (a) During pre-training, we use the same architecture as CroCo but replace the reconstruction objective with a co-visibility segmentation task, without masking, where each pixel in one view is classified as co-visible, occluded, or outside FOV with respect to the other view. (b) For fine-tuning on the relative pose regression task, we pool features from both views, process them through a shared MLP, and use separate heads for predicting translation and rotation.
Our architecture closely follows the design of CroCo, which consists of an encoder and a decoder. The encoder is a ViT [8] that processes the input images by dividing them into non-overlapping patches and encoding them into feature representations. The decoder is another transformer that combines information from both views, using cross-attention layers, to make predictions.
@@
As shown in Fig. 2 (a), given two images $I_{1}$ and $I_{2}$ of the same scene taken from different viewpoints, we first divide both into non-overlapping patches, denoted as tokens $p_{1}=\{p^{1}_{1},\ldots,p^{N_{1}}_{1}\}$ and $p_{2}=\{p^{1}_{2},\ldots,p^{N_{2}}_{2}\}$.
@@
$F_{1}=E_{\theta}(p_{1}),\quad F_{2}=E_{\theta}(p_{2})\>,$ (1)
@@
$O_{1}=D_{\phi}(F_{1},F_{2}),\quad O_{2}=D_{\phi}(F_{2},F_{1})\>,$ (2)
where $O_{1}$ and $O_{2}$ represent the output features from the decoder. The decoder contains cross-attention mechanisms that allow information exchange between features from both views.
3.2 Co-Visibility Segmentation Pre-Training
A key difference between our approach and CroCo lies in the pre-training objective. Instead of reconstructing masked pixels, we formulate the pre-training task as a co-visibility segmentation problem. Our goal is to predict for each pixel in each image whether it is:
@@ 3.2 Co-Visibility Segmentation Pre-Training
    Co-visible: the pixel corresponds to a 3D point that is also visible in the other image.
@@
$\hat{y}_{ikj}=\text{softmax}(W\cdot O_{ik}+b)_{j}\>,$ (3)
where $\hat{y}_{ikj}\in\mathbb{R}^{3}$ represents the predicted probabilities for the three co-visibility classes for pixel $j$ in patch $k$ of image $i$, and $W$ and $b$ are learnable parameters.
@@
$\mathcal{L}_{\text{ce}_{i}}=-\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\log(\hat{y}_{ij,c_{ij}})\>,$ (4)
where $c_{ij}$ is the ground-truth class in pixel $j$ of image $i$, and $N_{i}$ is the number of pixels in each image.
This formulation offers several advantages over the cross-view-completion-based approach. First, it allows us to use image pairs with any degree of overlap, as we explicitly model cases where pixels have no co-visible region in the other view. Second, it provides interpretable outputs that directly reveal the model’s understanding of scene geometry. Third, it better aligns with downstream tasks such as pose regression, where both images are fully visible.
@@ 3.3 Fine-Tuning for Relative Pose Regression
After pre-training, we fine-tune our model for the task of relative pose regression. We add a pose regression head on top of the pre-trained encoder-decoder architecture, while keeping the original co-visibility segmentation head. This design allows the model to leverage the geometric understanding acquired during pre-training.
The pose regression head consists of several components as shown in Fig. 2 (b). First, we apply global average pooling to the decoder outputs for both images:
$f_{1}=\text{GlobalAvgPool}(O_{1}),\quad f_{2}=\text{GlobalAvgPool}(O_{2})\>.$ (5)
@@
$f_{\text{shared}}=\text{MLP}([f_{1},f_{2}])\>.$ (6)
We then use separate heads for predicting the metric relative translation vector $t\in\mathbb{R}^{3}$ and the relative rotation represented as a quaternion $q\in\mathbb{R}^{4}$:
$t=\text{MLP}_{t}(f_{\text{shared}}),\quad q=\text{MLP}_{q}(f_{\text{shared}})\>.$ (7)
@@
    First, we freeze the pre-trained encoder, decoder and co-visibility segmentation head, and only train the pose regression head. During this phase, we use a homoscedastic loss [22] combining MSE losses for translation and quaternion predictions:
    $\mathcal{L}_{\text{pose}}=\frac{1}{2\sigma_{t}^{2}}\|t-\hat{t}\|^{2}+\frac{1}{2\sigma_{q}^{2}}\|q-\hat{q}\|^{2}+\log\sigma_{t}+\log\sigma_{q}\>,$ (8)
    where $t$ and $q$ are the ground-truth metric translation and quaternion respectively, and $\sigma_{t}$ and $\sigma_{q}$ are learnable parameters that automatically balance the two loss terms.
@@
    In the second phase, we unfreeze the backbone and the co-visibility segmentation head, and train the full network with a joint loss that combines the pose regression loss and the co-visibility segmentation loss:
    $\mathcal{L}_{\text{joint}}=\frac{1}{2\sigma_{\text{pose}}^{2}}\mathcal{L}_{\text{pose}}+\frac{1}{2\sigma_{\text{seg}}^{2}}\mathcal{L}_{\text{ce}}+\log\sigma_{\text{pose}}+\log\sigma_{\text{seg}}\>,$ (9)
    where $\sigma_{\text{pose}}$ and $\sigma_{\text{seg}}$ are learnable parameters that balance the two loss components.
This training strategy ensures that the model maintains its interpretable co-visibility segmentation capabilities while being optimized for the pose regression task.
4 Cub3: A Large-Scale Co-Visibility Dataset
To enable our co-visibility segmentation approach, we introduce Cub3, a large-scale dataset comprising 2.5 million image pairs with dense co-visibility annotations derived from the autonomous driving nuScenes dataset [2]. Cub3 will be made publicly available.

### v2

# Thibaut Loiseau1    Guillaume Bourmaud2    Vincent Lepetit1

@@ Abstract
Pre-training techniques have greatly advanced computer vision, with CroCo’s cross-view completion approach yielding impressive results in tasks like 3D reconstruction and pose regression. However, cross-view completion is ill-posed in non-covisible regions, limiting its effectiveness. We introduce Alligat0R, a novel pre-training approach that replaces cross-view learning with a covisibility segmentation task. Our method predicts whether each pixel in one image is covisible in the second image, occluded, or outside the field of view, making the pre-training effective in both covisible and non-covisible regions, and provides interpretable predictions. To support this, we present Cub3, a large-scale dataset with 5M image pairs and dense covisibility annotations derived from the nuScenes and ScanNet datasets. Cub3 includes diverse scenarios with varying degrees of overlap. The experiments show that our novel pre-training method Alligat0R significantly outperforms CroCo in relative pose regression. Code is available at https://github.com/thibautloiseau/alligat0r.
Figure 1: We introduce Alligat0R, a novel pretraining method for binocular vision. Alligat0R explicitly segments pixels as covisible, occluded, or outside field-of-view, overcoming the fundamental limitation of CroCo [weinzaepfel2022croco, weinzaepfel2023croco] which attempts to reconstruct potentially non-covisible regions.
@@ 1 Introduction
Pre-training techniques have revolutionized computer vision by enabling large models to learn rich representations [caron2021emerging, oquab2024dinov2, chen2020simple, he2022masked, grill2020bootstrap, bao2022beit, zhou2022image]. In 3D computer vision, CroCo [weinzaepfel2022croco, weinzaepfel2023croco] pioneered cross-view completion as a pretext task, where one view is partially masked and reconstructed using visible portions along with a second reference view of the same scene (see Figure˜1 left). This approach allows to learn powerful 3D features and has made possible impressive results on downstream tasks such as 3D reconstruction [wang2024dust3r, elflein2025light3r, yang2025fast3r, wang20243d], 3D pose regression [dong2024reloc3r], camera calibration [leroy2024grounding], 4D reconstruction [zhang2024monst3r, jin2024stereo4d, wang2025continuous] and Gaussian splatting [smart2024splatt3r, ye2024no].
Despite its success, the cross-view completion pretext task has a fundamental limitation: pretraining is only effective in covisible regions, as the task is ill-posed in non-covisible regions (see Figure˜1). As a consequence, Croco [weinzaepfel2022croco, weinzaepfel2023croco] relies on pairs with at least 50% of covisible regions.
In this paper, we present Alligat0R, a novel pre-training approach that offers an alternative to the cross-view completion objective through a covisibility segmentation task. Instead of reconstructing masked regions, our method explicitly predicts whether each pixel in one image is: (1) covisible in the second image, (2) occluded, or (3) outside the field of view (FOV). This formulation offers several advantages: it makes the pre-training effective in both covisible and non-covisible regions, it provides interpretable predictions that reveal the model’s geometric understanding, and it aligns more directly with the correspondence reasoning required in downstream binocular vision tasks.
To enable our approach, we introduce Cub3, a large-scale dataset comprising two sub datasets, each of 2.5 million image pairs with dense covisibility annotations derived from nuScenes [caesar2020nuscenes] and ScanNet [dai2017scannet] datasets. Cub3 includes challenging scenarios with varying degrees of overlap between views, providing a more diverse training signal than previous approaches.
One might notice that Alligat0R is not strictly self-supervised, unlike some earlier pre-training methods, as it relies on covisibility annotations. However, the same can be said for CroCo, which requires filtering image pairs to retain only those with at least 50% overlap. The creation of our covisibility annotations for nuScenes is more complex than for ScanNet, as the ground truth poses and depth maps are not available, however the process is fully automated. This process is very similar to the one used in RUBIK [loiseau2025rubik] and relies on registered video sequences and depth predictions.
@@
    We introduce covisibility segmentation as a novel pre-training objective for binocular vision tasks, replacing the cross-completion-based approach of prior work while maintaining the same network architecture.
@@
    We create and release Cub3, a large-scale dataset with dense covisibility annotations derived from nuScenes and ScanNet datasets.
@@
    We show that Alligat0R provides an effective alternative to CroCo pre-training when fine-tuned on the relative pose regression task, particularly for challenging scenarios with limited overlap between views.
@@
    We demonstrate the effectiveness of our pre-training and provide insights into the model’s cross-view reasoning capabilities through interpretable segmentation outputs.
Our experiments demonstrate that explicitly learning to understand covisibility relationships between image pairs leads to more robust and transferable features for relative pose regression compared to reconstruction-based approaches.
@@ 2 Related Work
To the best of our knowledge, CroCo [weinzaepfel2022croco] pioneered the extension of masked image modeling [hondru2024masked] to image pairs, introducing cross-view completion. CroCo v2 [weinzaepfel2023croco] applied this cross-view completion framework to large amounts of data. P-Match [zhu2023pmatch] introduced a variation where both images are partially masked to pre-train an image matching model. Another variant, masked appearance transfer, was proposed in [zhao2023representation] for object tracking, with a similar approach in [song2023compact]. Let us highlight that the foundational models DUSt3R [wang2024dust3r] and MASt3R [leroy2024grounding] are built on CroCo. While all these works rely on cross-view completion, we depart from this framework and propose a novel pretraining objective based on covisibility segmentation.
Considering covisibility across image pairs.
Ground-truth covisibility masks are widely used in image matching methods [potje2024xfeat, edstedt2024dedode, lindenberger2023lightglue, zhao2023aliked, sun2021loftr, wang2024efficient, edstedt2024roma, chen2022aspanformer, wang2022matchformer, tangquadtree, edstedt2023dkm, fan2023occ, gleize2023silk, kim2025learning, germainvisual, jiang2021cotr, tan2022eco, sarlin2020superglue] to compute overlaps and select training pairs. Several approaches [sarlin2020superglue, edstedt2024roma, truong2023pdc, edstedt2023dkm, germainvisual, bono2023end] also leverage these masks to learn matchability scores, which are used at test time to select correspondences. Additionally, [hutchcroft2022covispose] jointly learns covisibility and relative planar pose between 360 panoramas. To the best of our knowledge, our work is the first to learn visual representations from covisibility segmentation.
@@
Reloc3R [dong2024reloc3r] recently fine-tuned the foundational model DUSt3R [wang2024dust3r], which is based on CroCo, on large-scale data for relative pose regression. In our experiments, we adopt a similar strategy to evaluate our novel pretraining method against CroCo. However, we fine-tune Alligat0R for metric relative translation, whereas Reloc3R predicts only the direction of the relative translation.
@@ 3 Method
In this section, we describe our novel pre-training approach, Alligat0R, which replaces the cross-view completion objective used in CroCo [weinzaepfel2022croco, weinzaepfel2023croco] by a segmentation task. We first outline the overall architecture, which remains largely similar to CroCo, and then detail our covisibility segmentation pre-training objective. Finally, we explain how our model can be fine-tuned for downstream tasks such as relative pose regression.
@@ 3.1 Architecture Overview
Figure 2: Overview of Alligat0R. (a) During pre-training, we use the same architecture as CroCo but replace the reconstruction objective with a covisibility segmentation task, without masking, where each pixel in one view is classified as covisible, occluded, or outside FOV with respect to the other view. (b) For fine-tuning on the relative pose regression task, we pool features from both views, process them through a shared MLP, and use separate heads for predicting translation and rotation.
Our architecture closely follows the design of CroCo, which consists of an encoder and a decoder. The encoder is a ViT [dosovitskiy2020image] that processes the input images by dividing them into non-overlapping patches and encoding them into feature representations. The decoder is another transformer that combines information from both views, using cross-attention layers, to make predictions.
@@
As shown in Figure˜2 (a), given two images $\mathtt{I}_{1}$ and $\mathtt{I}_{2}$ of the same scene taken from different viewpoints, we first divide both into non-overlapping patches, denoted as tokens $\mathtt{P}_{1}=\{\mathtt{P}^{1}_{1},\ldots,\mathtt{P}^{N_{1}}_{1}\}$ and $\mathtt{P}_{2}=\{\mathtt{P}^{1}_{2},\ldots,\mathtt{P}^{N_{2}}_{2}\}$.
@@
$\mathtt{F}_{1}=E_{\theta}(\mathtt{P}_{1}),\quad\mathtt{F}_{2}=E_{\theta}(\mathtt{P}_{2})\>,$ (1)
@@
$\mathtt{O}_{1}=D_{\phi}(\mathtt{F}_{1},\mathtt{F}_{2}),\quad\mathtt{O}_{2}=D_{\phi}(\mathtt{F}_{2},\mathtt{F}_{1})\>,$ (2)
where $\mathtt{O}_{1}$ and $\mathtt{O}_{2}$ represent the output features from the decoder. The decoder contains cross-attention mechanisms that allow information exchange between features from both views.
3.2 Covisibility Segmentation Pre-Training
A key difference between our approach and CroCo lies in the pre-training objective. Instead of reconstructing masked pixels, we formulate the pre-training task as a covisibility segmentation problem. Our goal is to predict for each pixel in each image whether it is:
@@ 3.2 Covisibility Segmentation Pre-Training
    Covisible: the pixel corresponds to a 3D point that is also visible in the other image.
@@
$\hat{\mathtt{Y}}_{ikj}=\text{softmax}(\mathtt{W}\cdot\mathtt{O}_{ik}+\mathbf{b})_{j}\>,$ (3)
where $\hat{\mathtt{Y}}_{ikj}\in\mathbb{R}^{3}$ represents the predicted probabilities for the three covisibility classes for pixel $j$ in patch $k$ of image $i$, and $\mathtt{W}$ and $\mathbf{b}$ are learnable parameters.
@@
$\mathcal{L}_{\text{ce}_{i}}=-\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\ln(\hat{\mathtt{Y}}_{ij,c_{ij}})\>,$ (4)
where $\hat{\mathtt{Y}}_{ij,c_{ij}}$ is the predicted probability for the ground-truth class $c_{ij}$ in pixel $j$ of image $i$, and $N_{i}$ is the number of pixels in each image. The total cross-entropy loss is the sum over both images: $\mathcal{L}_{\text{ce}}=\mathcal{L}_{\text{ce}_{1}}+\mathcal{L}_{\text{ce}_{2}}$.
This formulation offers several advantages over the cross-view-completion-based approach. First, it makes the pre-training effective in both covisible and non-covisible regions, as we explicitly model cases where pixels have no covisible region in the other view. Second, it provides interpretable outputs that directly reveal the model’s understanding of scene geometry. Third, it better aligns with downstream tasks such as pose regression, where both images are fully visible.
@@ 3.3 Fine-Tuning for Relative Pose Regression
After pre-training, we fine-tune our model for the task of relative pose regression. We add a pose regression head on top of the pre-trained encoder-decoder architecture, while keeping the original covisibility segmentation head. This design allows the model to leverage the geometric understanding acquired during pre-training.
The pose regression head consists of several components as shown in Figure˜2 (b). First, we apply global average pooling to the decoder outputs for both images:
$\mathbf{f}_{1}=\text{GlobalAvgPool}(\mathtt{O}_{1}),\quad\mathbf{f}_{2}=\text{GlobalAvgPool}(\mathtt{O}_{2})\>.$ (5)
@@
$\mathbf{f}_{\text{shared}}=\text{MLP}([\mathbf{f}_{1},\mathbf{f}_{2}])\>.$ (6)
We then use separate heads for predicting the metric relative translation vector $\hat{\mathbf{t}}\in\mathbb{R}^{3}$ and the relative rotation represented as a quaternion $\hat{\mathbf{q}}\in\mathbb{R}^{4}$:
$\hat{\mathbf{t}}=\text{MLP}_{\mathbf{t}}(\mathbf{f}_{\text{shared}}),\quad\hat{\mathbf{q}}=\text{MLP}_{\mathbf{q}}(\mathbf{f}_{\text{shared}})\>.$ (7)
We use unit quaternions with L2 normalization, treating them as 4D vectors in the homoscedastic loss function.
@@
    First, we freeze the pre-trained encoder, decoder and covisibility segmentation head, and only train the pose regression head. During this phase, we use a homoscedastic loss [kendall2018multi] combining MSE losses for translation and quaternion predictions:
    $\mathcal{L}_{\text{pose}}=\frac{1}{2\sigma_{\mathbf{t}}^{2}}\|\mathbf{t}-\hat{\mathbf{t}}\|^{2}+\frac{1}{2\sigma_{\mathbf{q}}^{2}}\|\mathbf{q}-\hat{\mathbf{q}}\|^{2}+\log\sigma_{\mathbf{t}}+\log\sigma_{\mathbf{q}}\>,$ (8)
    where $\mathbf{t}$ and $\mathbf{q}$ are the ground-truth metric translation and quaternion respectively, and $\sigma_{\mathbf{t}}$ and $\sigma_{\mathbf{q}}$ are learnable parameters that automatically balance the two loss terms.
@@
    In the second phase, we unfreeze the backbone and the covisibility segmentation head, and train the full network with a joint loss that combines the pose regression loss and the covisibility segmentation loss:
    $\mathcal{L}_{\text{joint}}=\mathcal{L}_{\text{pose}}+\frac{1}{2\sigma_{\text{seg}}^{2}}\mathcal{L}_{\text{ce}}+\log\sigma_{\text{seg}}\>,$ (9)
    where $\sigma_{\text{seg}}$ is a learnable parameter.
This training strategy ensures that the model maintains its interpretable covisibility segmentation capabilities while being optimized for the pose regression task.
4 Cub3: A Large-Scale Covisibility Dataset
To enable our covisibility segmentation approach, we introduce Cub3, a large-scale dataset comprising two sub datasets each of 2.5 million image pairs with dense covisibility annotations derived from both the autonomous driving nuScenes dataset [caesar2020nuscenes], and the indoor ScanNet [dai2017scannet] dataset. Cub3 will be made publicly available.